# thanks: Corresponding authorthanks: Corresponding author

### v1

@@ Abstract
The direct detection of gravitational waves by LIGO has confirmed Einstein’s general relativity (GR) and sparked rapid growth in gravitational wave (GW) astronomy. However, subtle post-Newtonian (PN) deviations observed during the analysis of high signal-to-noise ratio events from the observational runs suggest that standard waveform templates, which assume strict adherence to GR, might overlook signals from alternative theories of gravity. Incorporating these exotic signals into traditional search algorithms is computationally infeasible due to the vast template space required. This paper introduces a deep learning framework for detecting exotic GW signals, leveraging neural networks trained on GR-based templates. Through their generalization ability, neural networks learn intricate features from the data, enabling the detection of signals that deviate from GR. We present the first study evaluating the capability of deep learning to detect beyond-GR signals, including a variety of PN orders. Our model achieves rapid and accurate identification of exotic GW signals across different luminosity distances, with performance comparable to GR-based detections. In particular, applying the model to the GW150914 event demonstrates excellent performance, highlighting the potential of AI-driven methods for detecting previously overlooked signals beyond GR. This work paves the way for new discoveries in gravitational wave astronomy, enabling the detection of signals that might escape traditional search pipelines.
@@ I Introduction
The detection of gravitational waves in 2015 marked the beginning of a new era in gravitational wave (GW) astronomy, providing a crucial test of general relativity (GR) under strong-field and high-energy conditions Abbott et al. (2016). To date, more than 90 GW events have been identified by the LIGO-Virgo-KAGRA collaboration Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2019); gwc ; Abbott et al. (2019a, 2021a, 2023a). These detections have significantly advanced our understanding of fundamental physics, cosmology, and astrophysics. In particular, they have provided insights into the formation and population distributions of compact objects Mandel and Broekgaarden (2022); van Son et al. (2022); Broekgaarden et al. (2021); Ezquiaga and Holz (2022) and enabled the measurement of cosmological parameters, such as the Hubble constant, using the standard siren method Abbott et al. (2017); Chen et al. (2018); Abbott et al. (2023b); Soares-Santos et al. (2019); Palmese et al. (2020); Abbott et al. (2021b); Holz and Hughes (2005); Dalal et al. (2006); Cutler and Holz (2009); Nakar and Sari (2010); Zhao et al. (2011); Camera and Nishizawa (2013); Vitale and Chen (2018); Bian et al. (2021); Cai and Yang (2017); Cai et al. (2018a); Cai and Yang (2018); Du et al. (2019); Cai et al. (2018b); Zhao et al. (2018); Zhang (2019); Yang et al. (2020, 2019); Gray et al. (2020); Bachega et al. (2020); Chang et al. (2019); Zhang et al. (2019); Mukherjee et al. (2021); He (2019); Zhao et al. (2020); Chen (2020); Chen et al. (2021); Mitra et al. (2021); Hogg et al. (2020); Nunes (2020); Borhanian et al. (2020); Jin et al. (2020); Yu et al. (2020); Wang et al. (2022); Qi et al. (2021); Jin et al. (2021); Cao et al. (2022); Leandro et al. (2022); Fu et al. (2021); Ye and Fishbach (2021); Zhu et al. (2022); de Souza et al. (2022); Jin et al. (2023a, 2022); Song et al. (2024); Jin et al. (2023b, 2024); Li et al. (2024a, b); Xiao et al. (2024).
@@
All currently reported gravitational wave events have been identified using both template-based Cannon et al. (2012); Privitera et al. (2014); Dal Canton et al. (2014); Usman et al. (2016); Messick et al. (2017); Sachdev et al. (2019) and non-template-based search methods Klimenko et al. (2016); Abbott et al. (2019b). Template-based searches rely on waveform templates constructed under the framework of general relativity (GR) to match signals in the detector data. In compact binary coalescence (CBC) searches, these templates represent theoretically precise models of the expected gravitational waveforms. The search algorithm evaluates the data to identify candidate triggers, assigning each a detection statistic to measure how well the observed data matches the template waveforms Abbott et al. (2018). However, when a signal diverges substantially from the expected GR waveforms—such as those predicted by beyond-GR (BGR) theories—template-based methods become significantly less effective. Given the wide variety of possible BGR models, it is computationally infeasible to create and deploy templates for every theoretical scenario. In contrast, non-template searches aim to detect excess power in the data across multiple detectors without assuming a specific waveform. These methods, in principle, can identify signals that deviate from standard GR predictions, including potential BGR events. However, the weaker modeling in non-template searches results in a higher susceptibility to noise artifacts that can mimic real signals. Consequently, these searches typically exhibit reduced sensitivity compared to template-based approaches, making it challenging to detect high-confidence signals. As a result, for most detected events, non-template searches have yielded lower sensitivity in comparison to template-based methods Abbott et al. (2019a, 2021a, 2023a). This makes it difficult to rely solely on non-template searches to identify and confirm BGR signals with high certainty. This dual approach underscores the trade-off between specificity and generality: while template-based searches offer high sensitivity for well-modeled signals, non-template searches provide broader coverage at the cost of reduced signal-to-noise discrimination. Both methods, however, are essential for advancing gravitational wave detection and testing new physics beyond GR.
@@
In recent years, deep learning, a subset of machine learning, has garnered significant attention for its successful application in gravitational wave (GW) detection Mukund et al. (2017); Gabbard et al. (2018); Fan et al. (2019); Krastev (2020); Wang et al. (2020); Krastev et al. (2021); Cabero et al. (2020); Wei et al. (2021); Schäfer et al. (2022); Verma et al. (2022); Moreno et al. (2022); Qiu et al. (2023); Zhang et al. (2022); Nousi et al. (2023); Ma et al. (2022); Schäfer et al. (2023); Wang et al. (2023); Yun et al. (2023); Ma et al. (2023); Trovato et al. (2024); Stergioulas (2024); McLeod et al. (2024); Wang et al. (2024); Sun et al. (2024); George and Huerta (2018); Xia et al. (2021); Jadhav et al. (2021); Zhao et al. (2023); Beveridge et al. (2023); Koloniari et al. (2024). By autonomously extracting patterns from data, deep learning has demonstrated the ability to analyze data with unknown characteristics, making it an invaluable tool for GW detection. Furthermore, its strong generalization capability—defined as the ability to accurately predict outcomes on unseen data Mohri et al. (2018)—is essential when working with real-world data, which is often more complex than simulated training datasets.
@@
This paper, for the first time, leverages the generalization capability of deep learning to search for BGR signals and explore the extent of this generalization. We employ BGR waveforms modeled using the parametrized post-Newtonian (pPN) formalism, which introduces deviations in the coefficients at various orders of the velocity parameter (v/c) within the phase expansion of the PN series. We employ a Residual Neural Network (ResNet) He et al. (2015), trained on GR-based signals, to investigate whether the network can generalize effectively to detect BGR signals, potentially revealing new physics beyond GR. The motivation behind this approach lies in the fact that accurately reconstructing the phase evolution of a signal provides the most sensitive method for detecting deviations from GR Li et al. (2012); Narola et al. (2023).
@@ II.1 Datasets preparation
We classify GW signals based on whether their templates include PN corrections. Notably, the BGR waveforms serve only for testing purposes and are not used during network training. GR templates are generated using PyCBC Nitz et al. (2023), simulating 1,000,000 one-second binary black hole (BBH) signals at a sampling rate of 4096 Hz, utilizing the IMRPhenomPv2 approximant Hannam et al. (2014); Khan et al. (2016); Husa et al. (2016). Of these, 800,000 waveforms are used for training, 100,000 for validation, and 100,000 for testing. The parameter distributions follow the settings used for GWTC-3 Abbott et al. (2023a), with minor modifications to the mass priors (see Table 1).
@@
Redshift Comoving (0, 2) -
Polarization angle Uniform (0, $\pi$) rad.
Spin magnitude Uniform (0, 0.998) -
Spin tilt Sine (0, $\pi$) rad.
@@
Declination angle Cosine ($-\pi/2$, $\pi/2$) rad.
Inclination angle Sine (0, $\pi$) rad.
The LIGO-Virgo collaboration has developed methods to quantify deviations from GR by parameterizing the inspiral, merger, and ringdown phases of detected events. Using the TIGER framework Li et al. (2012); Agathos et al. (2014); Meidam et al. (2018), they introduce fractional changes in the phase coefficients of the gravitational waveforms to model such deviations. Our work focuses on the inspiral phase, modeled using the TaylorF2 phase expansion $\Phi(f)$ Poisson and Will (1995); Buonanno et al. (2009):
@@
We use 10 days of publicly available background noise data from the Gravitational Wave Open Science Center (GWOSC) Abbott et al. (2023c), starting from GPS time 1130716817 (November 5th, 2015). From this dataset, we randomly sample 2,000,000 noise segments, each 2 seconds long. Half of these segments contain injected 1-second waveforms, with their peaks randomly distributed between 0.6 and 1.4 seconds. To prevent edge effects, any waveform exceeding the 2-second window is truncated. After estimating the noise power spectral density (PSD), we apply frequency-domain whitening before transforming the data back to the time domain. The middle 1 second of each segment is used for neural network analysis. This preparation ensures the network is trained on GR templates while the test phase explores its generalization ability with BGR signals.
@@ II.2 Deep learning model
For this study, we use a ResNet He et al. (2015), which incorporates skip connections to facilitate gradient flow and mitigate the vanishing gradient problem, allowing for efficient training of deeper networks. Deeper architectures, as shown in previous research, improve feature extraction and performance in complex tasks such as image recognition Simonyan and Zisserman (2014).
@@
Figure 3: Schematic of the ResNet network’s prediction process. The upper section shows data flow with sliding windows. The upward-pointing boundaries mark the start and end points of each window, while the downward-pointing arrows correspond to network outputs. Dashed windows indicate segments outside the network’s sensitive range. The lower section presents the sequence number of sliding windows on the horizontal axis and the corresponding network outputs on the vertical axis.
@@ II.3 Search strategy
During the training phase, three ResNet models are independently trained using data from different detectors: ResNetH for LIGO-Hanford, ResNetL for LIGO-Livingston, and ResNetHL, a dual-channel network trained on synchronized data from both detectors, as depicted in Fig. 2. ResNetHL processes the same time segments from both detectors simultaneously, outputting the probability that the two data segments contain signals originating from the same source.
@@
The sliding window mechanism ensures temporal overlap between consecutive segments, such that a signal merger time will ideally fall within two adjacent windows. If the difference between consecutive outputs is no more than 0.2, they are grouped as part of the same peak. Peaks containing 1 or 2 output values, each exceeding 0.4, are treated as signal alerts.
This peak detection approach is applied to all three networks. As shown in Fig. 2, the final prediction metric is obtained by summing the outputs of the three networks. This aggregated score ranges from 0 to 6, representing the combined confidence of the networks in detecting a signal within the analyzed data.
@@ III Results and Analysis
Deep learning models offer significant potential for detecting BGR signals, leveraging their ability to generalize from patterns observed in GR signals. These models, typically trained on simulated datasets, succeed with real observational data by identifying common features between the two, while mitigating differences. A central question explored here is whether a model trained solely on GR templates can generalize well enough to detect BGR signals and, if so, under what conditions this is possible.

### v2

@@ Abstract
The direct detection of gravitational waves by LIGO has confirmed Einstein’s general relativity (GR) and sparked rapid growth in gravitational wave (GW) astronomy. However, subtle post-Newtonian (PN) deviations observed during the analysis of high signal-to-noise ratio events from the observational runs suggest that standard waveform templates, which assume strict adherence to GR, might overlook signals from alternative theories of gravity. Incorporating these exotic signals into traditional search algorithms is computationally infeasible due to the vast template space required. This paper introduces a proof-of-principle deep learning framework for detecting exotic GW signals, leveraging neural networks trained on GR-based templates. Through their generalization ability, neural networks learn intricate features from the data, enabling the detection of signals that deviate from GR. We present the first study evaluating the capability of deep learning to detect beyond-GR signals, including a variety of PN orders. Our model achieves rapid and accurate identification of exotic GW signals across different luminosity distances, with performance comparable to GR-based detections. In particular, applying the model to the GW150914 event demonstrates excellent performance, highlighting the potential of AI-driven methods for detecting previously overlooked signals beyond GR. This work paves the way for new discoveries in gravitational wave astronomy, enabling the detection of signals that might escape traditional search pipelines.
@@ I Introduction
The detection of gravitational waves in 2015 marked the beginning of a new era in gravitational wave (GW) astronomy, providing a crucial test of general relativity (GR) under strong-field and high-energy conditions Abbott et al. (2016). To date, more than 90 GW events have been identified by the LIGO-Virgo-KAGRA collaboration Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2019); gwc ; Abbott et al. (2019a, 2021a, 2023a). These detections have significantly advanced our understanding of fundamental physics, cosmology, and astrophysics. In particular, they have provided insights into the formation and population distributions of compact objects Mandel and Broekgaarden (2022); van Son et al. (2022); Broekgaarden et al. (2021); Ezquiaga and Holz (2022) and enabled the measurement of cosmological parameters, such as the Hubble constant, using the standard siren method Abbott et al. (2017); Chen et al. (2018); Abbott et al. (2023b); Soares-Santos et al. (2019); Palmese et al. (2020); Abbott et al. (2021b); Holz and Hughes (2005); Dalal et al. (2006); Cutler and Holz (2009); Nakar and Sari (2010); Zhao et al. (2011); Camera and Nishizawa (2013); Vitale and Chen (2018); Bian et al. (2021); Cai and Yang (2017); Cai et al. (2018a); Cai and Yang (2018); Du et al. (2019); Cai et al. (2018b); Zhao et al. (2018); Zhang (2019); Yang et al. (2020, 2019); Gray et al. (2020); Bachega et al. (2020); Chang et al. (2019); Zhang et al. (2019); Mukherjee et al. (2021); He (2019); Zhao et al. (2020); Chen (2020); Chen et al. (2021a); Mitra et al. (2021); Hogg et al. (2020); Nunes (2020); Borhanian et al. (2020); Jin et al. (2020); Yu et al. (2020); Wang et al. (2022); Qi et al. (2021); Jin et al. (2021); Cao et al. (2022); Leandro et al. (2022); Fu et al. (2021); Ye and Fishbach (2021); Zhu et al. (2022); de Souza et al. (2022); Jin et al. (2023a, 2022); Song et al. (2024); Jin et al. (2023b, 2024); Li et al. (2024a, b); Xiao et al. (2024).
@@
All currently reported gravitational wave events have been identified using both template-based Cannon et al. (2012); Privitera et al. (2014); Dal Canton et al. (2014); Usman et al. (2016); Messick et al. (2017); Sachdev et al. (2019) and non-template-based search methods Klimenko et al. (2016); Abbott et al. (2019b). Template-based searches rely on waveform templates constructed under the framework of GR to match signals in the detector data. In compact binary coalescence (CBC) searches, these templates represent theoretically precise models of the expected gravitational waveforms. The search algorithm evaluates the data to identify candidate triggers, assigning each a detection statistic to measure how well the observed data matches the template waveforms Abbott et al. (2018). However, when a signal diverges substantially from the expected GR waveforms—such as those predicted by BGR theories—template-based methods become significantly less effective. Given the wide variety of possible BGR models, it is computationally infeasible to create and deploy templates for every theoretical scenario. In contrast, non-template searches aim to detect excess power in the data across multiple detectors without assuming a specific waveform. These methods, in principle, can identify signals that deviate from standard GR predictions, including potential BGR events. However, the weaker modeling in non-template searches results in a higher susceptibility to noise artifacts that can mimic real signals. Consequently, these searches typically exhibit reduced sensitivity compared to template-based approaches, making it challenging to detect high-confidence signals. As a result, for most detected events, non-template searches have yielded lower sensitivity in comparison to template-based methods Abbott et al. (2019a, 2021a, 2023a). This makes it difficult to rely solely on non-template searches to identify and confirm BGR signals with high certainty. This dual approach underscores the trade-off between specificity and generality: while template-based searches offer high sensitivity for well-modeled signals, non-template searches provide broader coverage at the cost of reduced signal-to-noise discrimination. Both methods, however, are essential for advancing gravitational wave detection and testing new physics beyond GR.
@@
In recent years, deep learning, a subset of machine learning, has garnered significant attention for its successful application in GW detection Mukund et al. (2017); Gabbard et al. (2018); Fan et al. (2019); Krastev (2020); Wang et al. (2020); Krastev et al. (2021); Cabero et al. (2020); Wei et al. (2021); Schäfer et al. (2022); Verma et al. (2022); Moreno et al. (2022); Qiu et al. (2023); Zhang et al. (2022); Nousi et al. (2023); Ma et al. (2022); Schäfer et al. (2023); Wang et al. (2023); Yun et al. (2023); Ma et al. (2023); Trovato et al. (2024); Stergioulas (2024); McLeod et al. (2024); Wang et al. (2024); Sun et al. (2024); George and Huerta (2018); Xia et al. (2021); Jadhav et al. (2021); Zhao et al. (2023); Beveridge et al. (2023); Koloniari et al. (2025). By autonomously extracting patterns from data, deep learning has demonstrated the ability to analyze data with unknown characteristics, making it an invaluable tool for GW detection. Furthermore, its strong generalization capability—defined as the ability to accurately predict outcomes on unseen data Mohri et al. (2018)—is essential when working with real-world data, which is often more complex than simulated training datasets.
@@
This paper, for the first time, leverages the generalization capability of deep learning to search for BGR signals and explore the extent of this generalization. We employ BGR waveforms modeled using the parametrized post-Newtonian (pPN) formalism, which introduces deviations in the coefficients at various orders of the velocity parameter (v/c) within the phase expansion of the PN series. We employ a Residual Neural Network (ResNet) He et al. (2015), trained on GR-based signals, to investigate whether the network can generalize effectively to detect BGR signals, potentially revealing new physics beyond GR. The motivation behind this approach lies in the fact that accurately reconstructing the phase evolution of a signal provides the most sensitive method for detecting deviations from GR Li et al. (2012); Narola et al. (2023). Although the deep learning architecture itself (i.e., ResNet) He et al. (2015) is established and has been successfully applied to gravitational wave detection in several previous studies Alhassan et al. (2024); Zelenka et al. (2024); Nagarajan and Messenger (2025), our novel contribution lies in the integration of three independent data streams within the search pipeline. This multi-stream approach not only enhances noise robustness but also addresses the shortcomings of conventional template-based methods for BGR detection.
@@ II.1 Datasets Preparation
Figure 1: Comparison of 1-second truncated waveforms at different match scores. From top to bottom are three sets of gravitational wave waveform comparisons, corresponding to match scores of 0.9, 0.6, and 0.2, respectively.
We classify GW signals based on whether their templates include PN corrections. Notably, the BGR waveforms serve only for testing purposes and are not used during network training. GR templates are generated using PyCBC Nitz et al. (2023), simulating 1,000,000 one-second binary black hole (BBH) signals at a sampling rate of 4096 Hz, utilizing the IMRPhenomPv2 approximant Hannam et al. (2014); Khan et al. (2016); Husa et al. (2016). Of these, 800,000 waveforms are used for training, 100,000 for validation, and 100,000 for testing. The parameter distributions follow the settings used for GWTC-3 Abbott et al. (2023a), with modifications to the mass priors to optimize neural network training. Specifically, we adopt uniform priors for the component masses ($m_{1}\in[5,100]M_{\odot}$ and $m_{2}\in[5,m_{1}]M_{\odot}$) rather than the astrophysically-motivated distributions used in GWTC-3. This modification addresses the oversampling of low-mass systems that occurs with standard GWTC-3 priors, ensuring balanced training coverage across the entire mass range and preventing network bias toward any particular mass region. These uniform priors are specifically chosen for machine learning training purposes and do not reflect assumptions about the true astrophysical mass distribution (see Table 1).
@@
Redshift Comoving (0, 2)
Polarization angle Uniform (0, $\pi$) rad
Spin magnitude Uniform (0, 0.998)
Spin tilt Sine (0, $\pi$) rad
@@
Declination angle Cosine ($-\pi/2$, $\pi/2$) rad
Inclination angle Sine (0, $\pi$) rad
The LIGO-Virgo collaboration has developed methods to quantify deviations from GR by parameterizing the inspiral, merger, and ringdown phases of detected events Agathos et al. (2014); Meidam et al. (2018); Abbott et al. (2021c). In the TIGER framework Li et al. (2012), they introduce fractional changes in the phase coefficients of the gravitational waveforms to model such deviations. Our work focuses on the inspiral phase, modeled using the TaylorF2 phase expansion $\Phi(f)$ Poisson and Will (1995); Buonanno et al. (2009):
@@
We use 10 days of publicly available background noise data from the Gravitational Wave Open Science Center (GWOSC) Abbott et al. (2023c), starting from GPS time 1130716817 (November 5th, 2015). From this dataset, we randomly sample 2,000,000 noise segments, each 2 seconds long. Half of these segments contain injected 1-second waveforms, with their peaks randomly distributed between 0.6 and 1.4 seconds. To prevent edge effects, any waveform exceeding the 2-second window is truncated. After estimating the noise power spectral density (PSD), we apply frequency-domain whitening before transforming the data back to the time domain. The middle 1 second of each segment is used for neural network analysis. This preparation ensures the network is trained on GR templates while the test phase explores its generalization ability with BGR signals. Notably, we deliberately train our network exclusively on GR templates, rather than including BGR waveforms in the training set. This choice, while potentially limiting the model’s specialization to specific BGR patterns, ensures broader generalization capability. Given the multitude of existing BGR theories and the uncertainty regarding which might accurately describe nature, training on GR templates alone helps maintain the model’s universality in detecting potential deviations from GR predictions.
Figure 2: Histograms showing the distribution of match values between pPN-modified and GR waveforms at different PN orders (0PN, 4PN, and 7PN). Solid lines represent distributions calculated using full waveforms, while dashed lines show distributions from signals truncated at 1 second. For each PN order, truncation shifts the distribution toward higher match values, but significant differences from GR remain detectable, with 0PN modifications showing the largest deviations (lowest match values).
PN effects primarily influence the inspiral phase of gravitational waveforms. Since our analysis applies a 1-second truncation to signals, we investigated whether this procedure significantly reduces the detectability of PN modifications.
Figure 1 displays truncated gravitational waveforms with varying match levels (0.9, 0.6, and 0.2). The comparison between GR waveforms (blue) and BGR predictions (orange) demonstrates that even with 1-second truncation, waveforms can exhibit substantial differences.
Figure 2 shows the distribution of match values between waveforms with pPN deviations and GR waveforms at different PN orders (0PN, 4PN, and 7PN). For each PN order, we compare the match distributions calculated using full signals (solid lines) versus signals truncated at 1 second (dashed lines). The histograms clearly demonstrate that while truncation shifts the distributions toward higher match values, significant differences between modified and GR waveforms persist across all PN orders. Notably, lower PN orders (particularly 0PN) show greater deviations from GR both before and after truncation, as indicated by their lower match values, while 4PN and 7PN modifications produce waveforms with progressively higher match values relative to GR predictions.
@@ II.2 Deep Learning Model
Deep learning is a subset of machine learning that specifically employs neural networks with multiple layers LeCun et al. (2015). While traditional machine learning methods often require manual feature engineering, deep learning models automatically learn hierarchical representations directly from raw data through their layered architecture Zeiler and Fergus (2014).
ResNet (Residual Network) introduces a key innovation to neural network design: skip connections (or shortcut connections) that allow information to bypass one or more layers He et al. (2015). These connections address the vanishing gradient problem that typically prevents the training of very deep networks. By enabling the direct flow of gradients through shortcuts, ResNet makes it feasible to train networks with hundreds or even thousands of layers, achieving superior performance on various pattern recognition tasks.
For this study, we use a ResNet. Deeper architectures, as shown in previous research, improve feature extraction and performance in complex tasks such as image recognition Simonyan and Zisserman (2014).
@@
Figure 5: Schematic of the ResNet network’s prediction process. The upper section shows data flow with sliding windows. The green boxes (labeled 1, 2, 3, … n, n+1, n+2) represent consecutive analysis windows, with red dashed lines marking the start and end positions of each window. The solid windows are within the network’s sensitive range, while dashed windows indicate segments outside this range. The lower section presents the sequence number of sliding windows on the horizontal axis and the corresponding network outputs (ranging from 0 to 1) on the vertical axis. The network outputs at adjacent windows can be summed when a signal spans multiple windows.
@@ II.3 Search Strategy
During the training phase, three ResNet models are independently trained using data from different detectors: ResNetH for LIGO-Hanford, ResNetL for LIGO-Livingston, and ResNetHL, a dual-channel network trained on synchronized data from both detectors, as depicted in Fig. 4. ResNetHL processes the same time segments from both detectors simultaneously, outputting the probability that the two data segments contain signals originating from the same source. The training and validation loss curves of the three models are shown in Appendix A.
@@
The sliding window mechanism ensures temporal overlap between consecutive segments, such that a signal merger time will ideally fall within two adjacent windows. If the difference between consecutive outputs is no more than 0.2, they are grouped as part of the same peak. Peaks with one or two output values exceeding 0.4 are considered signal alerts, a threshold optimized through extensive experimentation to balance sensitivity and the false alarm rate.
[…]
@@ III Results and Analysis
Deep learning models offer significant potential for detecting BGR signals, leveraging their ability to generalize from patterns observed in GR signals. These models, typically trained on simulated datasets, succeed with real observational data by identifying common features between the two, while mitigating differences. A central question explored here is whether a model trained solely on GR templates can generalize well enough to detect BGR signals and, if so, under what conditions this is possible. This approach represents a trade-off between biased generalization (which would result from training on specific BGR waveforms) and universal generalization (achieved by training only on GR templates), with our focus being on maintaining the latter to avoid premature assumptions about the nature of potential BGR signals.